In Desmos: Describe the transformations that occurred from the red parent graph f(x)= square root of x to the blue graph.


1
Expert's answer
2021-03-10T10:06:35-0500

vertical stretching is the stretching of the graph away from the x-axis

vertical compression (or shrinking) is the squeezing of the graph toward the x-axis.

• if > 1, the graph of k•f (x) is the graph of f (xvertically stretched by multiplying each of its y-coordinates by k.

• if 0 < k < (a fraction), the graph is f (xvertically shrunk (or compressed) by multiplying each of its y-coordinates by k.

• if k should be negative, the vertical stretch or shrink is followed by a reflection across the x-axis.

Notice that the "roots" on the graph stay in their same positions on the x-axis. The graph gets "taffy pulled" up and down from the locking root positions. The y-values change.

if our function is "f(x) = \sqrt{x}"









a) vertical compression/shrink by 1/2"\to"  "kf(x) = \large\frac{\sqrt{x}}{2}"

b) vertical stretch by 2 "\to kf(x) = 2\sqrt{x}"

c) Left 5 "\to f(x + k) = \sqrt{x+5}"

d) Right 5 "\to f(x - k) = \sqrt{x-5}"

e) Up 3 "\to f(x) + k = \sqrt{x} + 3"

f) Down 3 "\to f(x)-k = \sqrt{x} - 3"